Superman throws a 2400-N boulder at an adversary. What

horizontal force must Superman apply to the boulder to give it a

horizontal acceleration of 12.0 m/s2?


1
Expert's answer
2021-12-15T11:21:12-0500

"P = 240N \space\space\space\space a = 12\frac{m}{s^2}"

"F = ma = \frac{P}{g}a=\frac{240}{10}*12=288N"